OpenStudy (anonymous):

simplifly the sqaure roots 25/81

OpenStudy (anonymous):

\[\frac{5}{9}\]

OpenStudy (anonymous):

since \[\sqrt{\frac{a}{b}}=\frac{\sqrt{a}}{\sqrt{b}}\]

OpenStudy (anonymous):

what about 49/100

OpenStudy (anonymous):

\[\frac{7}{10}\]

OpenStudy (anonymous):

same reason

OpenStudy (anonymous):

7 over 10